OpenStudy (anonymous):

solve system of equations y=-2x+6 y=one halfx-9

OpenStudy (alexandervonhumboldt2):

y=-2x+6 y=1/2x-9 1/2x-9=-2x+6 5/2x=15 x=15*2/5=6